A current of 250. A flows for 24.0 hours at an anode where the reaction occurring is       Mn2+(aq) + 2H2O(l) → MnO2(s) + 4H+(aq) + 2e– What mass of MnO2 is deposited at this anode?

19.5 kg

4.87 kg

2.43 kg

12.9 kg

none of the above
Answer #1

Answer:

If the current is 250 A, it is also 250 C/s. This way we can find how many coulombs of electrons have passed through the anode if it was allowed to flow for 24.0 hrs.

24.0 hr * 60 min/ 1 hr * 60 s/ 1 min * 250 C/s = 2.16*10^7 C

Using Faraday's constant, we know that for every 96,485 C, there is 1 mol e-. So now we can see how many mols of electrons have been passed through.

2.16*10^7 C * 1 mol e-/96485 C = 223.87 mol e-

Using the balanced equation, since we know how many mols of electrons are present in the anode solution, we can see how many mols of MnO2 are deposited on the anode (by using a molar ratio):

Mn2+(aq) + 2H2O(l) → MnO2(s) + 4H+(aq) + 2e-

223.87 mol e- * 1 mol MnO2/ 2 mol e- = 111.93 mol MnO2

Using the molar mass of manganese dioxide (54.94 g/mol + (2* 15.9994 g/mol) = 86.94 g/mol), we can find the mass of MnO2.

111.93 mol MnO2 * 86.94 g MnO2/ 1 mol MnO2 = 9.73*10^3 g MnO2, or 9.73 kg MnO2has been desposited at the anode.